OpenStudy (anonymous):

rely need help factoring these. x:2x^2 + 7x + 5 = 0 x^2 - 4x - 5 = 0 9x^2 - 25 = 0 25x^2 - 1 = 0

OpenStudy (anonymous):

i dont have much time at all

OpenStudy (anonymous):

@4everaddicted2anime

OpenStudy (anonymous):

@rebeccaxhawaii

OpenStudy (anonymous):

pleas help gotta go in a minut

OpenStudy (anonymous):

Here is the first one.\[2 x^2+7 x+5=0\to (x+1) (2 x+5)=0 \]

OpenStudy (4everaddicted2anime):

Here is the second one: \[x^2-4x-5=0 \rightarrow (x-5)(x+1)=0\]

OpenStudy (mathmale):

Please, let's help nic123008 learn how to factor quadratics; that is not done by completing his/her work for him/her and giving away the answer. Thanks.